• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 00:04:51 on April 14, 2007

    Entry number 198944

    keyword=Shift Summary -- SWING April 13, 2007


    Shift Summary for April 13, 1600-2400


    Shift Crew:

    Shift Leader: Brad Sawatzky
    Target Operator: Mina Katramatou

    Kinematics:

    Q^2: 55 (GeV/c)^2
    Beam Energy: 0.9914 GeV
    Beam Current: 96 uA
    Left HRS: 0.59640 GeV/c @ 140.3 degrees
    Right HRS: 0.72620 GeV/c @ 14.9 degrees

    Summary:


    Time    Description
    1600    Started shift taking production data on He3
    1801    MCC removed beam to work on Hall B optics
    1807    Beam is back
    1922    Left Q1 tripped -- reset went smoothly
    1957    TS11 locked up at end of run 2426 -- reset from Hall A Tools GUI
    2030    Completed the shift checklist
    2320    No ADC data for LHRS S2 and Cer detectors
    .       - noticed in replay of run 2431)
    .       - Power cycled ROC3 and ROC4
    .       - Power cycled ROC2 (memory error when tried to CODA download)
    2340    Run 2433 quick replay looks good (ADCs are back)
    

    Run List:


    Run     Target      Events      Comments
    2418    He3         1.4M        Started during last shift
    2419    He3         1.4M
    2420    He3         1.4M
    2421    He3         1.4M
    2422    He3         0.35M       MCC took away beam to tweak Hall B optics
    2423    He3         1.4M
    2424    He3         1.4M
    2425    He3         0.4M        Left Q1 tripped at end of run
    2426    He3         1.4M        TS11 (left arm) froze at end of run
    2427    He3         1.4M
    2428    He3         1.4M
    2429    He3         1.4M
    2430    He3         1.4M
    2431    He3         1.4M        No S2-ADC or Cer-ADC in LHRS - JUNK?
    2432    He3         0.9M        No S2-ADC or Cer-ADC in LHRS - JUNK?
    2433    He3                     Handed of to next shift